OpenStudy (traylea):

Use point-slope form to write the equation of a line that has a slope of 2/3 and passes through (-3,-1). Write your final equation in slope-intercept form.

satellite73 (satellite73):

you know the point slope formula?

OpenStudy (traylea):

isnt it like y=mx+b

satellite73 (satellite73):

no

satellite73 (satellite73):

that is the "slope intercept" form \[\large y=\color{blue}mx+\color{red}b\] the slope is \(\color{blue}m\) and the y interecept is \(\color{red}b\)

satellite73 (satellite73):

what you want is the formula that uses a point \((\color{green}{x_1}, \color{magenta}{y_1})\) and a slope \(\color{blue}m\)

satellite73 (satellite73):

that formula is \[\large y-\color{magenta}{y_1}=\color{blue}m(x-\color{green}{x_1})\]

OpenStudy (traylea):

oh geez, I'm confused...can you walk me through?

satellite73 (satellite73):

point slope formula, what you use if you are given a point and a slope it is \[\large y-\color{magenta}{y_1}=\color{blue}m(x-\color{green}{x_1})\] your point is \( (\color{green}{-3},\color{magenta}{-1})\) and youir slope is \(\color{blue}{\frac{2}{3}}\)

satellite73 (satellite73):

put the numbers in the slots, then some algebra

OpenStudy (traylea):

So I would have to find the other coordinate right?

satellite73 (satellite73):

no not at all

satellite73 (satellite73):

there is no "other coordinate" but the three numbers in the place where they go i color coded them for you

satellite73 (satellite73):

*put

OpenStudy (traylea):

y - -1=2/3(x - -3)

satellite73 (satellite73):

yes that is a start

satellite73 (satellite73):

then change both \(--\) to a \(+\)

OpenStudy (traylea):

y+-1=2/3(x+-3) ???

satellite73 (satellite73):

no get rid of BOTH minus signs, make them one plus sign

OpenStudy (traylea):

so there will no longer be negatives

satellite73 (satellite73):

the same way that \[5-(-4)=5+4\]

satellite73 (satellite73):

zactly

OpenStudy (traylea):

sooo y+1=2/3(x+3)

satellite73 (satellite73):

yes

satellite73 (satellite73):

not done yet, but you did accomplish the first part "Use point-slope form to write the equation of a line that has a slope of 2/3 and passes through (-3,-1)."

satellite73 (satellite73):

next you have to "Write your final equation in slope-intercept form."

OpenStudy (traylea):

so now we do slope-intercept?

OpenStudy (traylea):

ok

satellite73 (satellite73):

two steps, always the same, to turn \[y-y_1=m(x-x_1)\] in to \[y=mx+b\]

satellite73 (satellite73):

first, the distributive property on the right \[y+1=\frac{2}{3}(x+3)\] distribute the \(\frac{2}{3}\) i.e. get rid of the parentheses on the right

OpenStudy (traylea):

y+1=2/3x+3

satellite73 (satellite73):

yes good

satellite73 (satellite73):

second step, subtract 1 from both sides to get y by itself on the left

satellite73 (satellite73):

oops nope sorry, not quite right

satellite73 (satellite73):

\[\frac{2}{3}\times 3=2\]

OpenStudy (traylea):

im confused lol

satellite73 (satellite73):

so \[y+1=\frac{2}{3}(x+3)\\ y+1=\frac{2}{3}x+2\]

OpenStudy (traylea):

-1 to only one side??

satellite73 (satellite73):

subtract 1 from both sides now

OpenStudy (traylea):

y+1-1=2/3x+2-1 y=2/3x+1

satellite73 (satellite73):

yup

OpenStudy (traylea):

is that it?

satellite73 (satellite73):

yes, now you are done it looks like \[y=mx+b\]

OpenStudy (traylea):

yay! thank you so much, you really helped me understand the concept.

satellite73 (satellite73):

yw good luck!